OpenStudy (anonymous):

Solve for a: 4a+7-3b=23

OpenStudy (arabpride):

4a + 7 - 3b = 23 -7 -7 4a - 3b = 16 +3b +3b 4a = 3b + 16 /4 /4 /4 a = 3/4b + 16/4 a = 3/4b + 4

OpenStudy (anonymous):

Move 7-3b to the RHS and then divide both sides by 4.